If Celsiusis -100 degree/boils what is fahrenheit?
Answer by dolly(163)

If C is celcius and F is farenheit then,
F= (9/5)*C + 32
= (9/5)*100 + 32
= 180 + 32
= 212
Thus 100 C = 212 F